# Campaign-spending driven polarization transition in a double-random field model of elections

### v2

@@ Abstract
A more recent class of opinion dynamics models builds on two sociological principles: homophily—the tendency to associate with similar others [24]—and social balance—the reduction of cognitive dissonance in triads [25, 26]. Extending the voter model, these approaches capture more realistic interactions. Their combined effects have been studied jointly [27, 28] and integrated into a unified framework [29], applied to group formation [30] and recently validated experimentally [31]. Many such models, including voter- and homophily-based ones, are inspired by the Ising model, long a versatile tool in interdisciplinary contexts [32]. In elections, it explained the universal scaling of vote distributions in proportional systems [33], predicted margins of victory from turnout [34], and modeled temporal effects through an external field [35].
@@
Despite its simplicity, the RFIM captures rich behavior such as quenched disorder and complex phase diagrams [37, 38, 39]. Variants with bimodal random fields, where the field takes two values [40, 41, 42], exhibit tricritical points marking the transition from second- to first-order phase changes [43, 44]. In our context, the RFIM represents a bipartisan electorate where each voter holds a binary preference. Voters occupy a social (friendship) network and are randomly assigned one of two field values: the sign encodes campaign affiliation, while the magnitude reflects campaign strength, with spending serving as a proxy. A schematic illustration is shown in End Matter, Fig. 3.
The aim of this paper is to study the interplay between homophily and campaign-following the emergence of campaign polarization, defined as the normalized difference between average opinions of voters exposed to each campaign. Low polarization indicates that both groups vote similarly, with decisions mainly shaped by homophily, whereas high polarization means groups align with campaigns and are less influenced by neighbors. We compute the phase diagram as a function of the “temperature,” representing susceptibility to opinion change, and the campaign spending of the two parties.
We focus on the effects of increasing campaign intensity. At low spending, opinions are shaped mainly by homophily, but as campaign influence grows, voter preferences are increasingly influenced by campaign messaging. Key questions are: when does campaign alignment outweigh homophilic similarity, how does polarization evolve at this point, and how does this transition affect social tension and election outcomes? A strength of our model is that it can be calibrated and tested on empirical data; we use particularly the US House elections between 1980 and 2020. Calibration allows us to infer the “temperature” and the critical spending threshold above which campaign polarization rises sharply, and to track how many races exceeded this threshold over four decades.
@@
In the End Matter, we present a similar analysis for races with $p\approx 0.6$, i.e., Republican-leaning districts ($p=0.6\pm 0.05$). Applied to this subset, the parameters shift slightly ($T=0.845$, $h^{*}\sim\mathdollar 2M$), but the qualitative behavior — hysteresis and emergence of polarization — remains unchanged.
@@
We shall also note that similar results were found in the context of polarization in the US Congress itself [64]. The study shows that the polarization in the US Congress increased after the 2010 Supreme Court approval of Super PACs, which enabled an increase in donor influence.
Elections are not the only context where homophily competes with external influence. Similar dynamics appear in marketing, where peer effects and advertising shape consumer behavior. A classic case is the Coca-Cola vs. Pepsi rivalry [65], showing how competing campaigns polarize preferences. Our framework illustrates how polarization generally arises from the tension between social ties and persuasive efforts.

### v3

# Empirical validation of the polarization transition in a double-random field model of elections

@@ Abstract
A more recent class of opinion dynamics models builds on two sociological principles: homophily—the tendency to associate with similar others [24]—and social balance—the reduction of cognitive dissonance in triads [25, 26]. Extending the voter model, these approaches capture more realistic interactions. Their combined effects have been studied jointly [27, 28] and integrated into a unified framework [29], applied to group formation [30], and recently validated experimentally [31]. Many such models, including voter- and homophily-based ones, are inspired by the Ising model, long a versatile tool in various interdisciplinary contexts [32], especially in opinion dynamics [33, 34, 35]. In elections, it explained the universal scaling of vote distributions in proportional systems [36], predicted margins of victory from turnout [37], and modeled temporal effects through an external field [38].
@@
Despite its simplicity, the RFIM captures rich behavior such as quenched disorder and complex phase diagrams [40, 41, 42]. Variants with bimodal random fields, where the field takes two values [43, 44, 45], exhibit tri-critical points marking the transition from second- to first-order phase changes [46, 47]. In the context of an election campaign, the RFIM represents a bipartisan electorate where each voter holds a binary preference. Voters occupy a social (friendship) network and are randomly assigned one of two field values: the sign encodes campaign affiliation, while the magnitude reflects campaign strength, with spending serving as a proxy. A schematic illustration is shown in End Matter, Fig. 3. RFIM approaches have long been used in sociophysics [48, 49], specifically to illustrate qualitative campaign effects [38], but have never been calibrated to an Ising-type model to with data on campaign spending.
The aim of this paper is to understand how the interplay between homophily and campaign-following leads to the emergence of campaign polarization in ways that can be calibrated to data. Campaign polarization is defined as the normalized difference between the average opinions of voters exposed to each campaign. Low polarization indicates that both groups vote similarly, with decisions mainly shaped by homophily, whereas high polarization means that groups align with campaigns and are less influenced by neighbors. We compute the phase diagram as a function of the “temperature,” representing susceptibility to opinion change, and the campaign spending of the two parties.
We focus on the effects of increasing campaign intensity. At low spending, opinions are shaped mainly by homophily, but as campaign influence grows, voter preferences are increasingly influenced by campaign messaging. Key questions are: when does campaign alignment outweigh homophilic similarity, how does polarization evolve at this point, and how does this transition affect social tension and election outcomes? A strength of our model is that it can be calibrated and tested on empirical data; we use particularly the US House elections between 1980 and 2020. Calibration allows us to infer the “temperature” and the critical spending threshold above which campaign polarization rises sharply, and to track how many races exceeded this threshold over four decades. To our knowledge, this is the first time where thermodynamic parameters and a critical spending threshold are directly extracted from historical election data and used to predict levels of polarization in society.
@@
In the End Matter, we present a similar analysis for races with $p\approx 0.6$, i.e., Republican-leaning races ($p=0.6\pm 0.05$) and $p\approx 0.4$, i.e., Democrat-leaning races. Applied to these two subsets, the parameters shift slightly ($T=0.845$ for Republican-leaning races, $T=0.865$ for Democrat-leaning races, $h^{*}\sim\mathdollar 2M$), but the qualitative behavior — hysteresis and emergence of polarization — remains unchanged.
@@
To further test the robustness of our approach, we present two additinal analyses in the Supplementary Material: in the first analysis, we divide the dataset in four decades and estimate the parameters for each decade separetely. We observe that while the temperature slightly decreases over time, the critical spending slightly increases. Second, we compare our approach to a machine learning approach based Support vector machine classification model. We show that the SVM can estimate the incumbency region (with slightly lower test accuracy), however, it cannot provide additional interpretion of the model, as the presence of polarization region.
@@
Note that similar results were found in the context of polarization in the US Congress itself [69]. The study shows that the polarization in the US Congress increased after the 2010 Supreme Court approval of Super PACs, which enabled an increase in donor influence. Also, other aspects such as social connectivity can increase polarization, not only in the context of elections [70, 71].
Elections are not the only context in which homophily competes with external influence. Similar dynamics appear in marketing, where peer effects and advertising shape consumer behavior. A classic case is the Coca-Cola vs. Pepsi rivalry [72], showing how competing campaigns polarize preferences. Our framework illustrates how polarization generally arises from the tension between social ties and persuasive efforts.
@@
Comparison of calibration of US House election data for different time periods.
To exemplify the robustness of the method and to investigate some aspects of the time-dependence of the thermodynamic quantities, we divide the data into four decades and estimate the model parameters separately. Since the number of data points for the original range of $p$ would be too small, we slightly increase it to $p=0.5\pm 0.1$. The fits are shown in Fig. 7, specifically, the incumbency regions are depicted in Fig. 8; the estimated parameters, together with the accuracy and average spending in the respective period, are summarized in Table 1. We observe that the temperature is decreasing slightly over time, while the critical spending is increasing. We also observe that the number of incumbents winning despite spending less decreases over time. The accuracy remains almost constant; its slightly smaller value (compared to the values in the main text) is due to the wider region of $p$.
Figure 7: Comparison of fitted parameters for US House elections in four decades.
Figure 8: Comparison of fitted parameters for US House elections in four decades corresponding to the previous plots, focused on incumbency regions
period # of races $T$ $h_{c}$ accuracy average spending (st.d.)
1980-1990 $581$ $0.925$ $\mathdollar 1.79M$ $0.8348$ $\mathdollar 759K$ ($\mathdollar 417K$)
1990-2000 $743$ $0.955$ $\mathdollar 1.43M$ $0.8318$ $\mathdollar 908K$ ($\mathdollar 501K$)
2000-2010 $515$ $0.900$ $\mathdollar 2.03M$ $0.8485$ $\mathdollar 1.50M$ ($\mathdollar 971K$)
2010-2020 $765$ $0.865$ $\mathdollar 2.31M$ $0.8353$ $\mathdollar 1.81M$ ($\mathdollar 1.39M$)
Table 1: Summary of estimated temperature $T$, critical spending $h_{c}$ and accuracy for each decade.
Comparison of calibration of US House election data with a support vector machine model. Finally, we compare our method with a standard machine learning classification method, particularly the support vector machine. Again, we focus on the close races, i.e., all races with $p=0.5\pm 0.05$. In order to utilize the natural symmetry of the system, i.e., $h^{+}=h^{-}$ leading to $m=0$, we transform the spending data into the following features:
$\sigma_{h}$ $=$ $\frac{{h^{+}}+{h^{-}}}{2}$
$\delta_{h}$ $=$ $h^{+}-h^{-}$
We then use the linear SVM on quadratic features, which is equivalent to the degree-2 polynomial SVM. Additionally, we add a binary feature indicating whether the winner was an incumbent. Since the training data is very unbalanced, we had to upsample the data to approximately equalize the number of instances: more spending by the winner and less spending by the incumbent. The trained accuracy is very high ($94\%$). By transforming the SVM back into the original space using
$h^{+}$ $=$ $\sigma_{h}+\delta_{h}/2$
$h^{-}$ $=$ $\sigma_{h}-\delta_{h}/2$
We transform the classifier into the original space. By measuring the intersection of the incumbency region with the diagonal ($h^{+}=h^{-}$), we obtain the estimate on the equivalent of the critical threshold, which is here $h^{*}=\mathdollar 2.68M$. This threshold is higher than predicted by the model in the main text. Furthermore, the accuracy on the whole dataset is lower (only $80\%$), possibly due to overfitting.
Figure 9: Application of SVM to estimate the incumbency region.